OpenStudy (anonymous):

You can place weights on both side of weighing balance and you need to measure all weights between 1 and 1000. For example if you have weights 1 and 3,now you can measure 1,3 and 4 like earlier case, and also you can measure 2,by placing 3 on one side and 1 on the side which contain the substance to be weighed. So question again is how many minimum weights and of what denominations you need to measure all weights from 1kg to 1000kg.

OpenStudy (wolf1728):

Seven weights 1 3 9 27 81 243 729

OpenStudy (wolf1728):

Or this is a LOT easier if you use powers of 2 1 2 4 8 16 32 64 128 256 512 This requires the weights in one pan only 1 = 1 2 = 2 3 = 2 + 1 4 = 4 5 = 4 + 1 6 = 4 + 2 7 = 4 + 2 + 1 8 = 8 9 = 8 + 1 etc.